OpenStudy (anonymous):

Solve the equation. 1/5y - 1 = 9 a) 45 b) 2 c) 40 d) 50 @KevinSlm , can you help?

OpenStudy (anonymous):

Add \(1\) to both the sides, in case if I am allowed to help..

OpenStudy (anonymous):

You are allowed. c:

OpenStudy (anonymous):

Then Add \(1\)..

OpenStudy (anonymous):

What will you get when you will add \(1\) both the sides?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

you get 1/5y = 10

OpenStudy (anonymous):

So would you do 1/5y =10/5 or no?

OpenStudy (anonymous):

This is right: \[\frac{1}{5}y =10\]

OpenStudy (anonymous):

Now, multiply by \(5\) both the sides and tell me what you got?

OpenStudy (anonymous):

The answer is 50?

OpenStudy (anonymous):

Yes it is..

OpenStudy (anonymous):

\[y =50\]

OpenStudy (anonymous):

Good..:)

OpenStudy (anonymous):

\[\dagger\]

OpenStudy (anonymous):

1/5y - 1 = 9 1/5y - 1 = 9 +1 1/5y - 1 = 10 1/5y - 1 = 10 *5 1/y = 50 1/y = 50 * 1 y = 50

OpenStudy (anonymous):

I wish I could give you another medal eta but there's a quota lol